OpenStudy (anonymous):

Critical numbers for cos9x=0 and sin9x=0?

myininaya (myininaya):

Do you mean find the critical numbers for f(x)=cos(9x) and g(x)=sin(9x)?

OpenStudy (anonymous):

Yep! @myininaya

myininaya (myininaya):

You need to start by differentiating f and g.

OpenStudy (anonymous):

Well the original equation was cos(9x)^2 and then I used the chain rule to get -18cos(9x)sin(9x) and now I'm trying to find the critical numbers. @myininaya

myininaya (myininaya):

So we have something like find the critical numbers for \[p(x)=(\cos(9x))^2\] Your p' is correct. \[p'(x)=-18\cos(9x)\sin(9x)\] Now set p'=0 Since we have a*b*c=0 this means a=0 or b=0 or c=0 -18=0 or cos(9x)=0 or sin(9x)=0 But -18 is never equal to 0 But we can solve the other two equations for x since cos and sin do be 0 sometimes actually infinitely many times do they be 0 So first cos(u)=0 For what values of u is cos 0?

myininaya (myininaya):

You might find the unit circle handy to look at.

myininaya (myininaya):

What is cos(pi/2)? cos(pi/2+pi)? cos(3pi/2+pi)? I hope you see a pattern here.

myininaya (myininaya):

Here I will show you how to solve this first equation but I want you to try to solve the sin(9x)=0. cos(9x)=0 cos(u)=0 when u=pi/2+n*pi where n is an integer But in place of 9x we put u so since u=9x then x=u/9 So if we want an answer in terms of x we say u=x/9=pi/2+n*pi And solve the equation for x by multiplying 9 on both sides like so x/9=pi/2+n*pi x=9pi/2+9n*pi

myininaya (myininaya):

You try solving sin(9x)=0

myininaya (myininaya):

lol googled this myself seen this question and seen a mistake :p and noticed a mistake I should have replaced u with 9x since u is 9x 9x=pi/2+npi divide both sides by 9 x=pi/18+npi/9 :p

myininaya (myininaya):

but also we didn't even need the derivative